# FinRobot: An Open-Source AI Agent Platform for Financial Applications using Large Language Models

### v1

@@ 2.4 AI4Finance Foundation and Open-Source Culture
The AI4Finance Foundation  [^1] https://www.ai4finance.org plays a crucial role in shaping the open-source culture within financial technology, driven by its mission to promote standardized practices and develop open-source resources. This commitment is aimed at benefiting both researchers and industry professionals, fostering a collaborative environment where industry-specific knowledge meets and innovation thrives. By leading initiatives in the application of AI technologies for financial services, AI4Finance not only accelerates technological advancements but also ensures these developments are accessible, transparent, and beneficial across the financial sector.
@@ 3 Overview of FinRobot: An Open-Source Platform for Financial Tasks
  • Multi-source LLM Foundation Models Layer: This foundational layer supports the plug-and-play functionality of various general and specialized LLMs. It forms the backbone of the platform, ensuring that all models are up-to-date, optimized, and consistently aligned with the latest advancements in financial technologies and data standards.
@@ 3.1.1 Multi-Agent Workflow
The Assistant in financial analysis focuses on initial data management tasks such as collecting, processing, and conducting preliminary analyses of financial data, and facilitates communication between agents. This role supports more specialized analysts by preparing datasets, maintaining databases, and executing basic financial computations and visualizations. This foundational work enables other analysts to focus on more complex and interpretive aspects of financial analysis.
@@ 3.3.2 Smart Scheduler Architecture
  • Task Manager: Manages and stores different general and fine-tuned LLMs based agents tailored for various financial tasks, updated periodically to ensure relevance and efficacy.
@@ 4 Financial Chain-of-Thought (CoT) Prompting
The ”chain-of-thought” prompting technique Wei et al. (2022); Xia et al. (2024) structures prompts to foster a step-by-step reasoning process within AI models, akin to human problem-solving strategies. This method significantly enhances performance on complex reasoning tasks such as mathematics and commonsense reasoning by encouraging models to articulate intermediate reasoning steps, leading to a final answer. This not only improves accuracy but also enhances the interpretability and transparency of the decision-making process.
4.1 Concept and Implementation
Financial CoT Prompting adapts the CoT technique for AI-driven financial analysis, integrating advanced cognitive processing techniques to refine decision-making capabilities. It stimulates the thought process of financial professionals to tackle complex financial issues by guiding AI models through a logical, sequential reasoning process. This approach breaks down complex financial scenarios into smaller, manageable components, analyzes each component, and synthesizes the findings to form conclusions or recommendations. It is particularly valuable for tasks requiring deep analytical thought, such as valuation, investment strategy formulation, market trend analysis, and risk assessment.
@@ 4.2 Advantages of Financial CoT Prompting
Despite their individual merits, a readily available tool that can intelligently perform industry- and company-specific analytics still does not exist in the way that parallels the work of an investment professional. Existing data providers such as Bloomberg, FactSet, CapitalIQ, and Refinitiv Eikon tend to rely on rudimentary, generic templates for data extraction, inadvertently overlooking key operational metrics and qualitative information intricately linked with the business essence. Manual analysis can be time-consuming and occasionally repetitive while existing automated tools often lack the capability to extract raw data from unstructured sources like annual reports. Although these platforms abound with vast amounts of information, the extraction of essential information and its subsequent analysis still require the expertise of investment professionals.
@@
With CoT Prompting, FinRobot establishes a unique niche among AI Agent Platforms by emulating human cognitive processes in financial analysis, thereby transcending the traditional reliance on mere numerical computations. Our methodology leverages LLMs to deconstruct and analyze the methodologies utilized by financial professionals when scrutinizing business data. This approach unveils that an effective analysis is intrinsically dependent on a comprehensive understanding of the business fundamentals of a company, as reflected through detailed accounting data, tailored industry-specific metrics, and qualitative information.
@@
This approach overcomes the above limitations because (1) it does not rely on a rigid template due to the generative nature of LLMs, (2) it avoids mis-classification because error-checking is embedded within layers of prompting, (3) it simplifies human labor by codifying investment logic within its reasoning, (4) it is able to extract relevant historical numbers and qualitative information from complex and unstructured data formats such as earning call transcripts, (5) it gives stronger grounds of explanation for the sources and derivations of each recorded and derived value given its multi-layered structure, and (6) it is highly adaptable and evolving as LLMs may be improved through supervised fine-tuning and instruction-tuning.
4.3 Financial CoT Prompting in Practice
Financial CoT Prompting represents a paradigm shift in the realm of security analysis, fundamentally altering conventional analytical tools across various domains:
  • Financial Analysis: FinRobot transcends mere data enumeration and ratio computation. It examines the interrelations among the three financial statements disclosed, discerns trends in historical financial data over time, and conducts comparative analyses across industry competitors. This comprehensive analytical approach is instrumental in gaining a profound understanding of the current business landscape of a company and delving into the significance and logical underpinnings of its financial metrics.
  • Business-Specific Analysis: Utilizing a bottom-up methodology to interpret publicly available information, FinRobot deconstructs multiple product lines, conducts trend analysis across various channels and regions, scrutinizes cost structures, evaluates dynamics within upstream and downstream supply chains, and assesses the conversion of research and development efforts. Additionally, it systematically comprehends a company’s business model, encompassing variations such as asset-light or asset-heavy models, subscription services, franchises, razor-and-blades strategies, network effects, and more. This meticulous analysis empowers FinRobot to offer a comprehensive understanding of the operational intricacies within a company, thereby facilitating informed decision-making and strategic planning.
  • Industry Analysis: Employing multi-layered reasoning, FinRobot dissects intricate datasets beyond a single company to unveil nuanced trends and underlying drivers across various metrics. Through this top-down approach to industry analysis, it not only enhances comprehension of industry dynamics and competitive landscapes but also anticipates how would these industry trends affects the company’s fundamentals. Consequently, investors and stakeholders are empowered to anticipate market shifts and make informed strategic decisions with foresight.
  • Investment Strategy: Building upon a comprehensive understanding of the financial statements, business model, and industry trends, FinRobot conducts evaluations encompassing a company’s future profitability, cash flow projections, and capital structure. Expanding on this foundation, FinRobot offers nuanced assessments spanning various investment horizons, structures, return criteria, and risk thresholds. Moreover, it provides strategic guidance regarding optimal investment timing and strategies, incorporating a range of financial instruments such as common equity investments, high-yield bonds, mezzanine financing, convertible bonds, derivatives, and other investment vehicles, all while delineating the uppermost limits of risk-return optimization.
@@ 5.1.3 Prompts
Aligning with the multi-task instruction tuning framework, the Market Forecaster follows a sophisticated prompt format. Specifically, the Market Forecaster gathers data of multifaceted company information, then it performs prompt engineering to format the instructive prompt with the structure of ‘Task Instruction + Company Information: Company Overview + Recent Stock Prices + Recent News + Latest Basic Financials’. Figure 5 shows an example of the Market Forecaster prompt template.
@@ 5.1.4 Sample Forecastings
The sample answers provided for two prominent stocks in the US and Chinese markets, namely AAPL and Kweichow Moutai, as depicted in Figures 6 and 7, respectively, illustrate the Market Forecaster’s adeptness in synthesizing and offering valuable insights gleaned from diverse information sources. Additionally, the Market Forecaster furnishes recommendations regarding the future trajectory of the stocks, underscoring its capability to provide actionable guidance based on the analyzed data.

### v2

@@ 2.4 AI4Finance Foundation and Open-Source Culture
The AI4Finance Foundation  [^1] https://www.ai4finance.org plays a crucial role in shaping the open-source culture within financial technology, driven by its mission to promote standardized practices and develop open-source resources. This commitment is aimed at benefiting both researchers and industry professionals, fostering a collaborative environment where industry-specific knowledge meets and innovation thrives. By leading initiatives in applying AI technologies for financial services, AI4Finance not only accelerates technological advancements but also ensures these developments are accessible, transparent, and beneficial across the financial sector.
@@ 3 Overview of FinRobot: An Open-Source Platform for Financial Tasks
  • Multi-source LLM Foundation Models Layer: This foundational layer supports the plug-and-play functionality of various general and specialized LLMs. It forms the platform’s backbone, ensuring that all models are up-to-date, optimized, and consistently aligned with the latest advancements in financial technologies and data standards.
@@ 3.1.1 Multi-Agent Workflow
The Assistant in financial analysis focuses on initial data management tasks such as collecting, processing, and conducting preliminary analyses of financial data and facilitates communication between agents. This role supports more specialized analysts by preparing datasets, maintaining databases, and executing basic financial computations and visualizations. This foundational work enables other analysts to focus on more complex and interpretive aspects of financial analysis.
@@ 3.3.2 Smart Scheduler Architecture
  • Task Manager: Manages and stores different general and fine-tuned LLMs-based agents tailored for various financial tasks, updated periodically to ensure relevance and efficacy.
@@ 4 Financial Chain-of-Thought (CoT) Prompting
4.1 Introduction
The chain-of-thought prompting technique Wei et al. (2022); Xia et al. (2024) structures prompts to foster a step-by-step reasoning process within AI models, akin to human problem-solving strategies. This method significantly enhances performance on complex reasoning tasks such as mathematics and commonsense reasoning by encouraging models to articulate intermediate reasoning steps, leading to a final answer. This method improves accuracy and enhances the interpretability and transparency of the decision-making process.
4.2 Concept of Financial CoT
Financial CoT Prompting adapts the CoT technique for AI-driven financial analysis, integrating advanced cognitive processing techniques to refine decision-making capabilities. It stimulates the thought process of financial professionals to tackle complex financial issues by guiding AI models through a logical, sequential reasoning process. This approach breaks down complex financial scenarios into smaller, manageable components, analyzes each component, and synthesizes the findings to form conclusions or recommendations. As noted by Kim et al. (2024), this method mimics human-like reasoning in the analysis chain and is particularly valuable for tasks requiring deep analytical thought, such as valuation, investment strategy formulation, market trend analysis, and risk assessment.
4.3 Financial CoT Implementation
Financial Chain-of-Thought (CoT) Prompting marks a transformative shift in security analysis, enhancing traditional analytical tools across various domains by leveraging the capabilities of LLMs such as GPT-4. Detailed case studies, such as those conducted by Kim et al. (2024), highlight this approach’s practical applications and benefits.
  • Financial Analysis: FinRobot transcends basic data enumeration and ratio computation. By employing LLMs for financial statement analysis, it conducts comparative analyses across industry competitors and against the company’s historical performance. This comprehensive approach provides profound insights into a company’s business landscape and the logical underpinnings of its financial metrics. Specifically, it normalizes financial ratios to identify anomalies, enhancing the accuracy of financial insights beyond human capabilities.
  • Business-Specific Analysis: Leveraging Retrieval-Augmented Generation (RAG), FinRobot revolutionizes the process of gathering data on a company’s products and services, complementing the information available in financial reports. By tapping into the vast resources of the web, FinRobot enriches its understanding of a company’s offerings, enabling a comprehensive analysis of product lines, trends across channels and regions, cost structures, supply chain dynamics, and R&D conversions.
  • Market Analysis: FinRobot abstracts real-world financial contexts, utilizing a combination of financial ratios, market data (including stock price trends and candlestick charts), market news sentiment, and alternative data sourced online. By employing LLMs, FinRobot simulates market participants’ decision-making processes in response to contextual changes. This allows for a comprehensive evaluation of a company’s stock price trajectory and valuation. FinRobot provides nuanced investment assessments, offering strategic guidance on optimal investment timing and suitable financial instruments, ensuring robust, well-informed investment strategies considering various horizons, structures, and risk thresholds.
  • Valuation Analysis: FinRobot integrates its comprehensive analysis of financial ratios, market data (including stock price trends and candlestick charts), market news sentiment, and alternative data collected online. FinRobot employs LLMs to evaluate a company’s stock price trajectory and valuation using the insights derived from these diverse sources. This approach enables FinRobot to provide nuanced investment assessments across various horizons, structures, and risk thresholds, offering strategic guidance on optimal investment timing and suitable financial instruments, thereby ensuring well-informed and robust investment strategies.
@@ 4.4 Advantages of Financial CoT Prompting
Despite their merits, a readily available tool that can intelligently perform industry- and company-specific analytics still needs to exist in a way that parallels the work of an investment professional. Existing data providers such as Bloomberg, FactSet, CapitalIQ, and Refinitiv Eikon rely on rudimentary, generic templates for data extraction, inadvertently overlooking key operational metrics and qualitative information intricately linked with the business essence. Manual analysis can be time-consuming and occasionally repetitive while existing automated tools can only sometimes extract raw data from unstructured sources like annual reports. Although these platforms abound with vast amounts of information, extracting essential information and its subsequent analysis still require the expertise of investment professionals.
@@
With CoT Prompting, FinRobot establishes a unique niche among AI Agent Platforms by emulating human cognitive processes in financial analysis, transcending traditional reliance on mere numerical computations. Our methodology leverages LLMs to deconstruct and analyze the methodologies utilized by financial professionals when scrutinizing business data. This approach unveils a practical analysis intrinsically dependent on a comprehensive understanding of a company’s business fundamentals, as reflected through detailed accounting data, tailored industry-specific metrics, and qualitative information.
@@
This approach overcomes the above limitations because: (1) It does not rely on a rigid template due to the generative nature of LLMs. (2) It avoids misclassification because error-checking is embedded within layers of prompting. (3) It simplifies human labor by codifying investment logic within its reasoning. (4) It can extract relevant historical numbers and qualitative information from complex and unstructured data formats such as earnings call transcripts. (5) Given its multi-layered structure, it provides stronger grounds for explaining the sources and derivations of each recorded and derived value. (6) It is highly adaptable and evolving as LLMs may be improved through supervised fine-tuning and instruction-tuning.
4.5 Market Simulation: Transcending the Analyst Mimicking
To expand the concept of market simulation and explore its possibilities, one could view the GPT’s ability to mimic human analytical processes as a gateway to simulating broader market dynamics. Market simulation transcends mere numerical analysis by incorporating human-like reasoning processes, embodying multiple personas and scenarios within a financial context. This simulation could encompass a variety of market participants whose decisions are influenced by an evolving set of data inputs—ranging from financial statements, reports, market data, and economic indicators to global news.
The core of this approach involves constructing a detailed simulation environment where these generative agents, representing different market players, interact Park et al. (2023). Each agent’s decision-making process would be guided by data-driven prompts, mimicking the cognitive steps a human analyst might take. This includes interpreting complex market signals and reacting to new information in real time, which advanced machine learning models and reinforcement learning techniques could facilitate.
By leveraging LLMs, one could theoretically encode these decision-making frameworks into the agents, enabling them to perform tasks that require a deep understanding of the market’s quantitative and qualitative aspects. This would simulate individual trading decisions and the broader market dynamics resulting from multiple decision-makers interactions. The potential to use this technology to create a virtual market environment where hypothetical scenarios could be tested and analyzed could provide invaluable insights into market behavior and strategy development.
A detailed framework for such a simulation could involve: (1) Defining the roles and behaviors of different market participants (e.g., institutional investors, retail investors, market makers). (2) Implementing a variety of data inputs and decision-making models for each participant. (3) Simulating market interactions over different time horizons and under varying conditions to assess potential outcomes and strategies. (4) Utilizing reinforcement learning to refine the agents’ strategies based on observed performance in the simulation.
As outlined by Park et al. (2023), this concept presents a promising avenue for future research and development in financial market analysis and strategy optimization.
@@ 5.1.3 Prompts
Aligning with the multi-task instruction tuning framework, the Market Forecaster follows a sophisticated prompt format. Specifically, the Market Forecaster gathers data of multifaceted company information, then it performs prompt engineering to format the instructive prompt with the structure of ‘Task Instruction & Company Information (Company Overview + Recent Stock Prices + Recent News + Latest Basic Financials)’. Figure 4 shows an example of the Market Forecaster prompt template.
@@ 5.1.4 Sample Forecastings
The sample answers provided for two prominent stocks in the US and Chinese markets, namely Nvidia and Kweichow Moutai, as depicted in Figures 5 and 6, respectively, illustrate the Market Forecaster’s adeptness in synthesizing and offering valuable insights gleaned from diverse information sources. Additionally, the Market Forecaster furnishes recommendations regarding the future trajectory of the stocks, underscoring its capability to provide actionable guidance based on the analyzed data.